OpenStudy (anonymous):

given the function described by g(x)=6x, find each of the following g(-12) g(16) g(30)

OpenStudy (anonymous):

It just means to plug in each of those values for x. g(-12) = 6(-12) = ? g(16) = 6(16) = ? etc

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

do i need to mutiple the equation 6*(-12 )

OpenStudy (anonymous):

Yes, that's what it means when you have numbers next to eachother like that with no operator between them. 6x = 6*x

OpenStudy (anonymous):

therefore g(-12)*6=-6

OpenStudy (anonymous):

g(-12) = 6*(-12) g(-12) = -36

OpenStudy (anonymous):

Ack!

OpenStudy (anonymous):

I need to wake up. g(-12) = 6*(-12) g(-12) = -72

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes